### v1

@@ Abstract
Strawberry harvesting robots faced persistent challenges such as low integration of visual perception, fruit-gripper misalignment, empty grasping, and strawberry slippage from the gripper due to insufficient gripping force, all of which compromised harvesting stability and efficiency in orchard environments. To overcome these issues, this paper proposed a visual fault diagnosis and self-recovery framework that integrated multi-task perception with corrective control strategies. At the core of this framework was SRR-Net, an end-to-end multi-task perception model that simultaneously performed strawberry detection, segmentation, and ripeness estimation, thereby unifying visual perception with fault diagnosis. Based on this integrated perception, a relative error compensation method based on the simultaneous target-gripper detection was designed to address positional misalignment, correcting deviations when error exceeded the tolerance threshold. To mitigate empty grasping and fruit-slippage faults, an early abort strategy was implemented. A micro-optical camera embedded in the end-effector provided real-time visual feedback, enabling grasp detection during the deflating stage and strawberry slip prediction during snap-off through MobileNet V3-Small classifier and a time-series LSTM classifier. Experiments demonstrated that SRR-Net maintained high perception accuracy. For detection, it achieved a precision of 0.895 and recall of 0.813 on strawberries, and 0.972/0.958 on hands. In segmentation, it yielded a precision of 0.887 and recall of 0.747 for strawberries, and 0.974/0.947 for hands. For ripeness estimation, SRR-Net attained a mean absolute error of 0.035, while simultaneously supporting multi-task perception and sustaining a competitive inference speed of 163.35 FPS. The compensation method reduced physical relative errors from 11.52 mm and 5.15 mm to 3.12 mm and 4.11 mm, below the tolerance threshold, while the early-abort strategy shortened execution time by 1.89 s and 0.3 s during deflating and snap-off stages. Overall, the proposed framework enhanced both the robustness and efficiency of strawberry harvesting robots, providing a practical and reliable solution for autonomous fruit harvesting.
@@ 1 Introduction
In the HarvestFlex system (https://xiong-lab.cn/), as shown in Fig. 1 (a) and (b), a RealSense D455F depth camera (Intel, Realsense, USA) captured real-time RGB-D images, which were processed by YOLOv8/11 for strawberry detection and segmentation. A checkerboard-based calibration enabled eye-to-hand coordinate transformation, while polynomial regression compensated spatial errors. For sequence planning, a minimal-height sorting algorithm determined harvesting order, and a point-to-point speed pattern-based method ensured efficient move the end-effector. The end-effector was a flexible pneumatic gripper that inflated and deflated to gently envelop and detach strawberries. The complete harvesting workflow consisted of six steps: inflating and approaching, swallowing, deflating, snap off, descending and placing, in Fig. 2. However, several issues emerged with our HarvestFlex robot that severely compromised its ability to harvest stably and continuously. These primary challenges were: (1) Low integration of visual perception tasks – the limited fusion of detection, segmentation, and ripeness estimation led to increased model complexity and reduced inference speed, thereby constraining overall system efficiency. (2) Positional inaccuracy between the end-effector and the target fruit – although the eye-to-hand approach offered a stable field of view for dynamic fruit tracking, it remained highly sensitive to hand-eye calibration errors. Even minor deviations could lead to misalignment, empty grasps, unsuccessful detachments, or incorrect placements, thereby disrupting continuous operation. (3) Inadequate gripping and strawberry slippage – although the pneumatic soft gripper reduced fruit damage, its high compliance makes the integration of reliable tactile or slippage sensors challenging. Consequently, slippage often went undetected, resulting in wasted cycle time and, in some cases, unintended fruit release during the snap-off stage. This problem was further compounded by the immaturity of current tactile sensing technologies, which remained expensive, unstable, and impractical for large-scale deployment in orchard environments. Moreover, existing tactile sensor technologies remained immature, with solutions that are typically expensive, unstable, and challenging to deploy reliably in orchard environments.
@@
To overcome these limitations, this paper focused on a vision-based fault diagnosis framework for HarvestFlex, thereby avoiding the need for additional multi-modal sensors and the complexities of data alignment. The objective was to address software- and control-related faults that compromised harvesting stability and efficiency. To this end, an end-to-end multi-task perception method SRR-Net [SUN2026111169] was introduced and integrated visual perception with fault object detection while maintaining a lightweight model architecture and low computational load. The method comprised three primary subtasks: object detection, instance segmentation, and ripeness estimation. To align the gripper with the harvesting point, a relative error compensation method based on the simultaneous target-gripper detection was implemented in the coordinate frame of robot arm once the gripper reached the position beneath the target. Then, a corrected harvesting point was generated, guiding the gripper to move beneath the target for a second time to ensure accurate alignment. Furthermore, an early abort strategy was introduced to improve the reliability of the harvesting process. Specifically, a micro-optical camera embedded at the base of the gripper continuously monitors the presence and stability of strawberries. In the deflating stage, the MobileNet V3-Small [9008835] verified whether the fruit had been successfully grasped, while in the snap-off stage, an LSTM classifier estimated the probability of the strawberry slippage from the gripper. If no strawberry was detected, an early abort signal was triggered to terminate the current cycle. This integrated perception–action framework enabled prompt detection of empty grasps or slippage, allowing the robotic arm to respond rapidly and maintain stable operation. The main contributions of this paper were as follows.
@@ 2 Dataset Benchmark
In this paper, FaultData, GraspData and SnapData were constructed to support the multi-task vision perception task, grasp adjustment task and enable the time-series LSTM classifier task. All data were collected using the HarvestFlex strawberry-harvesting robot, which includes two robotic arms [chen2024design], two Realsense D455F cameras, and two controllable light sources, as shown in Fig. 1(a–b). The cameras captured RGB-D images at a resolution of 640×480 pixels, with the distance between the lens and the strawberries ranging from 30 to 90 cm. Data were collected in both natural and controlled lighting environments to enhance the robot’s visual adaptability to complex and dynamic orchard conditions.
@@
FaultData For the tasks of detection, segmentation and ripeness estimation of strawberry and gripper, diverse strawberry formations-such as isolated, overlapping, and occluded fruits-as well as the complete operational sequence of the end-effector during picking images were collected. To ensure robust and continuous operation for multi-task vision perception task, the dataset also incorporated a variety of lighting conditions, weather scenarios, and nighttime environments, laying the foundation for stable, autonomous, all-weather performance. In addition, high-resolution images were captured using a Redmi Note 13 Pro smartphone and an Orbbec Gemini Pro RGB-D camera (Orbbec, Gemini Pro, China) providing enhanced texture and color information. The entire dataset was collected at the Cuihu Factory in Beijing, China, and featured the Fragaria × ananassa ‘Kaorino’ cultivar, as shown in Fig. 1(c). All images were annotated with the polygonal outlines of strawberries and the end-effector using Labelme [torralba2010labelme]. Each object instance was labeled as strawberry, hand or table, with the ripeness of strawberries assigned based on [SUN2026111169]. The ripeness score was defined within the range [0,1.1]. The label format for each instance followed the structure: <<cls>> <<ripeness>> <<boundaries>>, where cls represented the object class. The dataset was split into training and validation subsets, consisting of 2954 and 779 images, respectively. Note that the ripeness attribute applied only to strawberries. To ensure a consistent label format, the table and hand class was assigned alignment flags of 2 and 3.
@@
GraspData Using a miniature camera, image data were collected during the deflation stage of the end-effector. A total of 594 images were acquired. In the training and validation sets, the numbers of images with strawberries and without strawberries during the deflating stage are 197/202 and 93/102, respectively.
@@
SnapData For the time-series prediction task during the strawberry snap-off stage, all images were captured using a micro-optical camera mounted on the end-effector based on artificial strawberry instances. In each frame, key visual features—including the normalized strawberry area, normalized gripper area, normalized background area, and the width and height of the strawberry—were extracted using a fine-tuned SRR-Net. These features were structured in the format: <<strawberry_area>>, <<gripper_area>>, <<background_area>>, <<w>>, <<h>>, <<label>>, where label denotes the presence (0) or absence (1) of the strawberry within the gripper. A sliding window of 10 consecutive frames was used to form each input sequence. For each sequence, a label was assigned based on the next 3 frames immediately following the sequence. This label indicated whether the strawberry was at risk of slipping from the gripper. This structure enabled the training and validation of a binary LSTM classifier to estimate the likelihood of strawberry slippage. The training and validation sets comprise 341 and 147 samples, respectively.
@@ 3.1 Overview
The HarvestFlex robot has successfully achieved strawberry harvesting in orchard environments. However, several challenges persisted that hindered harvesting performance during continuous operations. The core challenges included: (a) limited integration of visual perception; (b) grasping misalignment or failure resulting from positional offsets; and (c) reliance on human intervention to determine success or failure during the picking process, owing to the absence of automated evaluation and recovery mechanisms, as shown in Fig. 3.
@@
To address the aforementioned issues, a vision-based harvest fault diagnosis and self-recovery architecture was developed. This architecture integrated three key components: an end-to-end multi-task perception network, a relative error compensation method, an early abort strategy based on the empty grasp adjustment and strawberry slip prediction. Specifically, the end-to-end multi-task perception network, SRR-Net, simultaneously performed detection, segmentation, and ripeness estimation. Detection, segmentation, and ripeness estimation were integrated into an end-to-end method. A vision-based positional error compensation method was implemented to correct the $x$- and $y$-axis offsets between the picking point and the gripper center, thereby enhancing localization accuracy. To detect unsuccessful grasps, a micro-optical camera was embedded within the end-effector to monitor strawberry retention during the deflating stage and predict potential slippage risks during the snap-off stage. A MobileNet V3-Small classifier was applied to determinate the whether strawberry are present at the end-effector during the deflating stage. An LSTM-based visual time-series prediction model provided real-time forecasts of slippage, enabled a secondary harvesting attempt, the triggering of an early abort signal, or the continuation of subsequent actions.
@@ 3.2 End-to-End Multi-Task Perception Method: Detection, Segmentation and Ripeness Estimation
An end-to-end multi-task perception method, SRR-Net, was introduced to simultaneously perform detection, segmentation, and ripeness estimation of both strawberries and gripper. This method integrated low- and high-level feature representations with ripeness and biological attributes to comprehensively perceive strawberries and reduce errors in ripeness estimation. The method employed a lightweight network with shared weights. For strawberries, the perception task distinguished between ripe and unripe fruits via post-processing classification. For the robotic gripper, relative error between the gripper and the picking point was calculated.
@@ 3.3 Simultaneous Target-Gripper Detection for Relative Error Compensation
In the HarvestFlex arm’s workflow, a relative error compensation method was incorporated into the inflating and approaching and swallowing stages of the picking process, in Fig. 5.
When the gripper approached the strawberry picking point, the 3D coordinates of both the end-effector and the picking point in the robot arm coordinate system were obtained. Specifically, using SRR-Net, the 2D image coordinates of the strawberry and the end-effector were extracted in the camera coordinate system. By combining these with the corresponding depth information, their 3D positions in the camera coordinate system were reconstructed and then transformed into the robot arm coordinate system through hand–eye calibration. Let $(x_{p},y_{p},z_{p})$ and $(x_{e},y_{e},z_{e})$ denote the coordinates of the picking point and the end-effector in the robot arm frame, respectively. The positional error between the end-effector and the picking point was defined as a 3D offset ($\Delta x$, $\Delta y$, $\Delta z$) , defined as
@@
Since the snap-off stage necessitated bending the strawberry stem, the vertical position of gripper during the swallowing stage was adjusted to exceed that of the picking point. To simplify computation and reduce processing overhead, only the $x$- and $y$-axis errors were considered, while $z$-axis errors were ignored. Given an error tolerance threshold $T$, when the end-effector reached a position beneath the picking point and the absolute error exceeded $T$, the compensated picking point coordinates $(x_{c},y_{c},z_{c})$ were computed as:
$x_{c}=\begin{cases}x_{p}+\Delta x&\text{if }|\Delta x|>T\\ x_{p}&\text{otherwise}\end{cases}$ (3)
$y_{c}=\begin{cases}y_{p}+\Delta y&\text{if }|\Delta y|>T\\ y_{p}&\text{otherwise}\end{cases}$ (4)
@@
Here, $T$ denoted the maximum tolerance positional error for each axis, set to 10. The robotic arm was then moved to the compensated point for strawberry grasping.

### v2

@@ Abstract
Strawberry harvesting robots faced persistent challenges such as low integration of visual perception, fruit-gripper misalignment, empty grasping/misgrasp, and strawberry slippage from the gripper due to insufficient gripping force, all of which compromised harvesting stability and efficiency in orchard environments. To overcome these issues, this paper proposed a visual fault diagnosis and self-recovery framework that integrated multi-task perception with corrective control strategies. At the core of this framework was SRR-Net, an end-to-end multi-task perception model that simultaneously performed strawberry detection, segmentation, and ripeness estimation, thereby unifying visual perception with fault diagnosis. Based on this integrated perception, a relative error compensation method based on the simultaneous target-gripper detection was designed to address positional misalignment, correcting deviations when error exceeded the tolerance threshold. To mitigate empty grasping/misgrasp and fruit-slippage faults, an early abort strategy was implemented. A micro-optical camera embedded in the end-effector provided real-time visual feedback, enabling grasp classification during the deflating stage and strawberry slip prediction during snap-off through MobileNet V3-Small classifier and a time-series LSTM classifier. Experiments demonstrated that SRR-Net maintained high perception accuracy. For detection, it achieved a precision of 0.895 and recall of 0.813 on strawberries, and 0.972/0.958 on hands. In segmentation, it yielded a precision of 0.887 and recall of 0.747 for strawberries, and 0.974/0.947 for hands. For ripeness estimation, SRR-Net attained a mean absolute error of 0.035, while simultaneously supporting multi-task perception and sustaining a competitive inference speed of 163.35 FPS. The proposed compensation mechanism reduced the relative errors from 11.52 mm and 5.15 mm to 3.12 mm and 4.11 mm, while maintaining an execution time below 1 s. In addition, the grasp adjustment module responded within 0.5 s, leading to a reduction of about 0.5 s in the deflating stage. During the snap-off stage, when strawberry slipping was predicted, re-inflation followed by a secondary snap-off required 1.88 $\pm$ 0.01 s, achieving a recovery success rate of 81.25%. In cases where slipped was predicted and the operation was terminated, the total execution time was 1.44 $\pm$0.07 s, with slipped status success rate of 88.89%. Compared with the original cycle time of 11.21 $\pm$ 0.15 s, grasp adjustment reduced the cycle time to 4.56 $\pm$ 0.04 s, representing a reduction of approximately 6.7 s, while strawberry slip prediction reduced it to 7.08 $\pm$ 0.10 s, corresponding to a reduction of approximately 4 s. Overall, the proposed framework enhanced both the robustness and efficiency of strawberry harvesting robots, providing a practical and reliable solution for autonomous fruit harvesting.
@@ 1 Introduction
In the HarvestFlex system (https://xiong-lab.cn/), as shown in Fig. 1 (a) and (b), a RealSense D455F depth camera (Intel, Realsense, USA) captured real-time RGB-D images, which were processed by YOLOv11 for strawberry detection and segmentation. A checkerboard-based calibration enabled eye-to-hand coordinate transformation, while polynomial regression compensated spatial errors. For sequence planning, a minimal-height sorting algorithm determined harvesting order, and a point-to-point speed pattern-based method ensured efficient move the end-effector. The end-effector was a flexible pneumatic gripper [11231358] that inflated and deflated to gently envelop and detach strawberries. The complete harvesting workflow consisted of six steps: inflating and approaching, swallowing, deflating, snap off, descending and placing, in Fig. 2. However, several issues emerged with our HarvestFlex robot that severely compromised its ability to harvest stably and continuously. These primary challenges were: (1) Low integration of visual perception tasks – the limited fusion of detection, segmentation, and ripeness estimation led to increased model complexity and reduced inference speed, thereby constraining overall system efficiency. (2) Positional inaccuracy between the end-effector and the target fruit – although the eye-to-hand approach offered a stable field of view for dynamic fruit tracking, it remained highly sensitive to hand-eye calibration errors. Even minor deviations could lead to misalignment, empty grasps, unsuccessful detachments, or incorrect placements, thereby disrupting continuous operation. (3) Inadequate gripping and strawberry slippage – although the pneumatic soft gripper reduced fruit damage, its high compliance makes the integration of reliable tactile or slippage sensors challenging. Consequently, slippage often went undetected, resulting in wasted cycle time and, in some cases, unintended fruit release during the snap-off stage. This problem was further compounded by the immaturity of current tactile sensing technologies, which remained expensive, unstable, and impractical for large-scale deployment in orchard environments. Moreover, existing tactile sensor technologies remained immature, with solutions that are typically expensive, unstable, and challenging to deploy reliably in orchard environments.
@@
To overcome these limitations, this paper focused on a vision-based fault diagnosis framework for HarvestFlex, thereby avoiding the need for additional multi-modal sensors and the complexities of data alignment. The objective was to address software- and control-related faults that compromised harvesting stability and efficiency. To this end, an end-to-end multi-task perception method SRR-Net [SUN2026111169] was introduced and integrated visual perception with fault object detection while maintaining a lightweight model architecture and low computational load. The method comprised three primary subtasks: object detection, instance segmentation, and ripeness estimation. To align the gripper with the harvesting point, a relative error compensation method based on the simultaneous target-gripper detection was implemented in the coordinate frame of robot arm once the gripper reached the position beneath the target. Then, a corrected harvesting point was generated, guiding the gripper to move beneath the target for a second time to ensure accurate alignment. Furthermore, an early abort strategy was introduced to improve the reliability of the harvesting process. Specifically, a micro-optical camera embedded at the base of the gripper continuously monitors the presence and stability of strawberries. In the deflating stage, the MobileNet V3-Small [9008835] verified whether the fruit had been successfully grasped, while in the snap-off stage, an LSTM classifier estimated the probability of the strawberry slippage from the gripper. If the strawberry was slipping, re-inflation and snap-off actions were triggered. If it had already slipped, an early abort signal was issued to terminate the current cycle. This integrated perception–action framework enabled prompt diagnosis of empty grasps/misgrasp, and slippage, allowing the robotic arm to respond rapidly and maintain stable operation. The main contributions of this paper were as follows.
@@ 2 Dataset Benchmark
In this paper, FaultData, GraspData, SnapData and SlipData were constructed to support the multi-task vision perception task, grasp adjustment task and enable the time-series LSTM classifier task. All data were collected using the HarvestFlex strawberry-harvesting robot, which includes two robotic arms [chen2024design], two Realsense D455F cameras, as shown in Fig. 1(a–b). The cameras captured RGB-D images at a resolution of 640×480 pixels, with the distance between the lens and the strawberries ranging from 30 to 90 cm. Data were collected in both natural and controlled lighting environments to enhance the robot’s visual adaptability to complex and dynamic orchard conditions.
@@
FaultData For the tasks of detection, segmentation, and ripeness estimation of strawberry and gripper, diverse strawberry formations-such as isolated, overlapping, and occluded fruits-as well as the complete operational sequence of the end-effector during picking images were collected. To ensure robust and continuous operation for multi-task vision perception task, the dataset also incorporated a variety of lighting conditions, weather scenarios, and nighttime environments, laying the foundation for stable, autonomous, all-weather performance. LightStrawberry from [SUN2026111169] also was applied into FaultData. In addition, high-resolution images were captured using a Redmi Note 13 Pro smartphone and an Orbbec Gemini Pro RGB-D camera (Orbbec, Gemini Pro, China) providing enhanced texture and color information. The entire dataset was collected at the Cuihu Factory in Beijing, China, and featured the Fragaria × ananassa ‘Kaorino’ cultivar, as shown in Fig. 1(c). All images were annotated with the polygonal outlines of strawberries and the end-effector using Labelme [torralba2010labelme]. Strawberry with the ripeness of strawberries assigned based on [SUN2026111169] and hand were labeled. The ripeness score was defined within the range [0, 1.1]. The label format for each instance followed the structure: <<cls>> <<ripeness>> <<boundaries>>, where cls represented the object class. The dataset was split into training and validation subsets, consisting of 2954 and 779 images, respectively. Note that the ripeness attribute applied only to strawberries. To ensure a consistent label format, the hand class was assigned alignment flags of 2.
@@
GraspData Using a miniature camera, image data were collected during the deflation stage of the end-effector. The dataset comprised three classes: 389 images of ripe strawberries held in the gripper (label 0), 346 images of empty grasps with no strawberry (label 1), and 388 images of unripe strawberries held in the gripper (label 2). Following a 7:3 stratified split, the training set contained 272, 242, and 272 images for the three classes respectively, while the validation set contained 117, 104, and 116 images.
[…]
@@
SlipData In each frame, key visual features—including the normalized strawberry area, normalized gripper area, normalized background area, and the width, height and center point of the strawberry—were extracted using a fine-tuned SRR-Net based on the SnapData. These features were structured in the format: <<strawberry_area>>, <<gripper_area>>, <<background_area>>, <w><w>, <h><h>, <x><x>, <y><y>, <<label>>. This label indicated whether the strawberry was at risk of slipping from the gripper, with three possible statuses: normal (0), slipping (1), and slipped (2). A sliding window of 5 consecutive frames was used to form each input sequence. For each sequence, a label was assigned based on the next 3 frames immediately following the sequence. This structure enabled the training and validation of a LSTM classifier to estimate the likelihood of strawberry slippage. The sample counts for labels 0, 1, and 2 were 719, 157, and 1962, respectively. Due to the brief duration of the slipping status during the snap-off stage, there is severe class imbalance, particularly for samples labeled as 1 (slipping). To mitigate the issue of sample imbalance, an oversampling method was introduced. Then, each category was split into training and validation sets at a 7:3 ratio. The images from frame 2 to frame 13 were presented, with the frame number and status displayed in the upper left corner of each subfigure in left of Fig. 3. On the right side were the variation curves of the seven feature components throughout the snap-off process, along with their corresponding labels. The horizontal axis represented the frame number, the vertical axis represented the feature values of the corresponding features, and the blue line indicates the labels for the corresponding frames.
Figure 3: Strawberry slip process and status
@@ 3.1 Overview
The HarvestFlex robot has successfully achieved strawberry harvesting in orchard environments. However, several challenges persisted that hindered harvesting performance during continuous operations. The core challenges included: (a) limited integration of visual perception; (b) grasping misalignment or failure resulting from positional offsets; and (c) reliance on human intervention to determine success or failure during the picking process, owing to the absence of automated evaluation and recovery mechanisms, as shown in Fig. 4. A video demo was available at https://1drv.ms/v/c/e68b602748988e28/IQDbhbliJ1nPTqooSIfJg0S0AeDAfkL3JQGtFZfRcWEQImA?e=ISC7Mm.
@@
To address the aforementioned issues, a vision-based harvest fault diagnosis and self-recovery architecture was developed. This architecture integrated three key components: an end-to-end multi-task perception network, a relative error compensation method, an early abort strategy based on the empty grasp/misgrasp adjustment and strawberry slip prediction. Specifically, the end-to-end multi-task perception network, SRR-Net, simultaneously performed detection, segmentation, and ripeness estimation. Detection, segmentation, and ripeness estimation were integrated into an end-to-end method. A vision-based positional error compensation method was implemented to correct the $x$- and $y$-axis offsets between the picking point and the gripper center, thereby enhancing localization accuracy. To detect unsuccessful grasps, a micro-optical camera was embedded within the end-effector to monitor strawberry retention during the deflating stage and predict potential slippage risks during the snap-off stage. A MobileNet V3-Small classifier was applied to determine the whether ripe strawberry are present at the end-effector during the deflating stage. An LSTM-based visual time-series prediction model provided real-time forecasts of slippage, enabled a secondary harvesting attempt, the triggering of an early abort signal, or the continuation of subsequent actions.
@@ 3.2 End-to-End Multi-Task Perception Method: Detection, Segmentation and Ripeness Estimation
An end-to-end multi-task perception method, SRR-Net, was introduced to simultaneously perform detection, segmentation, and ripeness estimation of both strawberries and gripper. This method integrated low- and high-level feature representations with ripeness and biological attributes to comprehensively perceive strawberries and reduce errors in ripeness estimation. The method employed a lightweight network with shared weights. For strawberries, the perception task distinguished between ripe and unripe fruits by post-processing classification. For the robotic gripper, relative error between the gripper and the picking point was calculated.
@@ 3.3 Simultaneous Target-Gripper Detection for Relative Error Compensation
In the HarvestFlex arm’s workflow, a relative error compensation method was incorporated into the inflating and approaching and swallowing stages of the picking process, in Fig. 6. When the gripper approached the strawberry picking point, the 3D coordinates of both the end-effector and the picking point were obtained in the robot arm coordinate system. Specifically, using SRR-Net, the 2D image coordinates of the strawberry and the end-effector were extracted in the camera coordinate system. By combining these with the corresponding depth information, their 3D positions in the camera coordinate system were reconstructed and then transformed into the robot arm coordinate system through hand–eye calibration. Let $(x_{p},y_{p},z_{p})$ and $(x_{e},y_{e},z_{e})$ denote the coordinates of the picking point and the end-effector in the robot arm frame, respectively. The positional error between the end-effector and the picking point was defined as a 3D offset ($\Delta x$, $\Delta y$, $\Delta z$) , defined as
@@
Since the snap-off stage required bending the strawberry stem, the vertical position of gripper during the swallowing stage was adjusted to exceed that of the picking point. To simplify computation and reduce processing overhead, only the $x$- and $y$-axis errors were considered, while $z$-axis errors were ignored. Given an error tolerance threshold $T$, when the end-effector reached a position beneath the picking point and the absolute error exceeded $T$, the compensated picking point coordinates $(x_{c},y_{c},z_{c})$ were computed as:
$x_{c}=\begin{cases}x_{p}+k_{x}\Delta x&\text{if }|\Delta x|>T\\ x_{p}&\text{otherwise}\end{cases}$ (3)
$y_{c}=\begin{cases}y_{p}+k_{y}\Delta y&\text{if }|\Delta y|>T\\ y_{p}&\text{otherwise}\end{cases}$ (4)
@@
Here, $T$ denoted the maximum tolerance of positional error for each axis, set to 10 mm. $k_{x}$ and $k_{y}$ referred to the scale factors of error. The robotic arm was then moved to the compensated point for strawberry grasping.